A 76.00 pound flask of mercury costs $132.00.The density of mercury is 13.534 g/cm3.

Find the price of one cubic inch of mercury by calculating intermediate values.

What is the price of one pound of mercury?

$

What is the price of 1 g of mercury?

$

What is the price of 1 cm3 of mercury?

$

What is the price of 1 in3 of mercury?

$

It takes 5.000 in3 of mercury to make one manometer. Find the price of the mercury used to make 19 manometers by first calculating the cost of mercury for one manometer.

What is the price of mercury used to make one manometer?

$

What is the price of mercury used to make 19 manometers?
Answer #1

Answer:-

(I)-

wt. of mercury = 76.00 pound

cost of 76.00 pound of mercury = 132.00 $

As we know that

cost of 76.00 pound of mercury = 132.00 $

then

cost of 1.0 pound of mercury = 132.00 / 76.00 $

cost of 1.0 pound of mercury = 1.74 $

So  price of one pound of mercury is 1.74 $ ( i.e the answer)

(II)-

wt. of mercury = 76.00 pound

Since

1.00 pound = 453.592 g

So

wt. of mercury = 76.00 pound = 76.00 453.592 g = 34472.992‬ g

therefore

cost of 34472.992‬ g of mercury = 132.00 $

As we know that

cost of 34472.992‬ g of mercury = 132.00 $

then

cost of 1.0 g of mercury = 132.00 / 34472.992‬ $

cost of 1.0 g of mercury = 0.0038 $

So  price of 1 g of mercury is 0.0038 $ ( i.e the answer)

(III)-

wt. of mercury = 76.00 pound

Since

1.00 pound = 453.592 g

So

wt. of mercury = 76.00 pound = 76.00 453.592 g = 34472.992‬ g

density of mercury = 13.534 g/cm3

As we know that

mass = volume density

mass or wt. of mercury = volume of mercury   density of mercury

34472.992‬ g = volume of mercury   13.534 g/cm3

volume of mercury = 34472.992‬ g / 13.534 g/cm3

volume of mercury =  2547.140 cm3

Since we know that

cost of 76.00 pound of mercury = cost of 34472.992‬ g of mercury = cost of 2547.140 cm3  of mercury = 132.00 $

So

cost of 2547.140 cm3  of mercury = 132.00 $

then

cost of 1.0 cm3  of mercury = 132.00 / 2547.140 $

cost of 1.0 cm3  of mercury = 0.052 $

So  price of 1 cm3 of mercury is 0.052 $ ( i.e the answer)

(IV)-

wt. of mercury = 76.00 pound = 76.00 453.592 g = 34472.992‬ g

density of mercury = 13.534 g/cm3

As we know that

mass = volume density

mass or wt. of mercury = volume of mercury   density of mercury

34472.992‬ g = volume of mercury   13.534 g/cm3

volume of mercury = 34472.992‬ g / 13.534 g/cm3

volume of mercury =  2547.140 cm3

since

1 cm3 = 0.0610237 in3

then

volume of mercury =  2547.140 cm3 = 2547.140 0.0610237 in3

volume of mercury = 155.436 in3

So

cost of 2547.140 cm3  of mercury = cost of 155.436 in3 of mercury = 132.00 $

therefore

cost of 155.436 in3 of mercury = 132.00 $

then

cost of 1 in3 of mercury = 132.00 / 155.436  $

cost of 1 in3 of mercury = 0.85  $

So  price of 1 in3 of mercury is 0.85 $ ( i.e the answer).

(V)-

As we know that

volume of mercury used to make 1 manometer = 5.00 in3

Since we know that

price of 1 in3 of mercury = 0.85 $

then

price of 5.00 in3  of mercury = 5.00 0.85 $

price of 5.00 in3  of mercury = 4.25 $

price of mercury used to make one manometer = 4.25 $

So price of mercury used to make one manometer is 4.25 $ ( i.e the answer).

(VI)-

As we know that

price of mercury used to make one manometer = 4.25 $

then

price of mercury used to make 19 manometer = 19 4.25 $

price of mercury used to make 19 manometer = 80.75‬ $

So price of mercury used to make 19 manometers is 80.75‬ $ ( i.e the answer).